Jallouk Corporation has two different bonds currently outstanding. Bond M has a face value of $20,000 and matures In 20 years. The bond makes no payments for the first six years, then pays $1,300 every six months over the subsequent eight years, and finally pays $1.600 every six months over the last six years. Bond N also has a face value of $20,000 and a maturity of 20 years. It makes no coupon payments over the life of the bond. The required return on both these bonds is 8 percent compounded semiannually. 


What is the current price of Bond M and Bond N? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) 



Current price
Bond M
Bond N
Answer #1
BOND M
Particulars Year Amount
Value of 6 years Coupon payment and maturity value at the end of year 14 14 $27,508.06
Value of Coupon Payment made during the year 6.5 to 14 years and present value compounded at the end of 14 year 6 $29,834.76
Present Value of Bond 0 $18,634.70
So, the Present Value of Bond M is $18,634.70
Bond N
Present Value $4,165.78
Thus,the present value of bond N is $4,165.78

Working Note:-

Present Value Calculation Table
BOND M
Particulars Year Amount
Value of 6 years Coupon payment and maturity value at the end of year 14 14 =-PV(4%,12,1600,20000,0) $27,508.06
Value of Coupon Payment made during the year 6.5 to 14 years and present value compounded at the end of 14 year 6 =-PV(4%,16,1300,27508.06,0) $29,834.76
Present Value of Bond 0 =-PV(4%,12,0,29834.76,0) $18,634.70
So, the Present Value of Bond M is $18,634.70
Present Value of Bond N
Bond N
Present Value =-PV(4%,40,0,20000,0) $4,165.78
Thus,the present value of bond N is $4,165.78
Note:- Return 8 % compounded semi annually so to convert Componded per semi annual = 8/2 =4%

Similary year will be double so for 6 year per semiannually year will be 6*2=12years